OpenStudy (anonymous):

I have to integrate this question and show my substitutions. My question is integral 2/(t^3-4t) dt to start this question do I move the bottom to the top? and make it to the ^-1?

OpenStudy (agreene):

I would start by factoring out the 2, and then use partial fractions...

OpenStudy (anonymous):

there is only a 2 on the top. I can factor a t from the bottom

OpenStudy (agreene):

the 2 is a constant.

OpenStudy (anonymous):

I was thinking of putting integral 2(t^3-4t)^-1 dt? I really struggle with this stuff

OpenStudy (agreene):

\[\int \frac{2}{t^3-4t}dt=2\int\frac{1}{t^3-4t}dt\]

OpenStudy (anonymous):

oh yeah forgot about the constant.

OpenStudy (agreene):

\[2\int (-\frac{1}{4t}+\frac{1}{8(t+2)}+\frac{1}{8(t-2)})dt\] from partial fractions, from there I would split into the 3 different fractional integrands and go from there--they look fairly easy. Of course, if you haven't learned partial fractions--probably not the best idea to use it, lol... but, thats where my mind goes when I see this type of question.

OpenStudy (anonymous):

no we have not learned partial fractions that is next on the agenda.

OpenStudy (anonymous):

all the examples we have are involving NON fractions thats why having trouble.

OpenStudy (agreene):

hrm, honestly I'm not sure how to do this without using partial fractions....

OpenStudy (anonymous):

hmmm... maybe I will read ahead then and see how to do partial fractions then. You are obviously very intelligent. thanks :)